12 men can complete a work in 4 days. how many men will be required in order to complete 8 times work in half times?
Answer by ankor@dixie-net.com(22740)

12 men can complete a work in 4 days.
Find how many man-days
12 * 4 = 48 man-days to complete one job
:
Assume this means
how many men will be required in order to complete 8 times the work, in half the time?
Half the time = 2 days, let m the no. of men to do this
2m = 8*48
2m = 384
m = 384/2
m = 192 men to do 8 times the work in 2 days